OpenStudy (anonymous):

A system of equations is shown below: 6x-5y=5 and 3x/5y=4 I If any answer you if you ask me a question because I'm not on my computer. Sorry but pleaseeeeeeeee help me!!!!

OpenStudy (anonymous):

Can someone please help ? I understand how to do the problem but I can't figure it out

OpenStudy (anonymous):

the second equation...is that right 3x/5y=41

OpenStudy (anonymous):

No 4

OpenStudy (anonymous):

3x/5y=4 right

OpenStudy (anonymous):

by grahing you get x=20/21 and y=1/7

OpenStudy (ranga):

6x-5y=5 --- (1) 3x/5y=4 --- (2) multiply both sides of (2) by 5y 3x = 20y --- (3) 3x - 20y = 0 multiply by 2 6x - 40y = 0 subtract from equation 1: 0 +35y = 5 y = 5/35 = 1/7 From (3) x = 20/3 * y = 20/3 * 1/7 = 20/21 x = 20/21, y = 1/7